OpenStudy (tparker22):

Write the point-slope form of the line that passes through (1, -5) and is perpendicular to a line with a slope of 1.

OpenStudy (leenathan):

is there a graph?

OpenStudy (tparker22):

No, they just gave me that equation and they want me to write out how I worked it. I'm totally stuck

OpenStudy (leenathan):

oh hmmm wait ill try not sure if ican do its maybe @jabez177 @pooja195 or @Austin.L

OpenStudy (tparker22):

Ok, thanks!

jabez177 (jabez177):

Point slope form Do you remember the form?

OpenStudy (tparker22):

y - y1 = m(x - x1)

OpenStudy (leenathan):

i would say (-1, 5) jab should know if im right

jabez177 (jabez177):

Ohhhh!! This form? Eeks... This is my least favorite one... Practice makes perfect. :)

OpenStudy (leenathan):

XD

OpenStudy (tparker22):

Ugh, I hate it so much! >_< My teacher isn't really that big of a help either. I'm like, "Dude, I'm going for broadcast journalism. I don't need this stupid point slope form crap"

jabez177 (jabez177):

Alright so first off. We have to plug in the numbers

pooja195 (pooja195):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @tparker22 y - y1 = m(x - x1) \(\color{#0cbb34}{\text{End of Quote}}\) You have this info down just plug in your given numbers

pooja195 (pooja195):

m=slope

jabez177 (jabez177):

y – y1 = m(x – x1) So we just put in 5 where the Y1 is and the -1 where the x1 is

jabez177 (jabez177):

You get the picture?

OpenStudy (tparker22):

Ok, so would that be y - y6 = m(x - x-1)

pooja195 (pooja195):

Take out the y you are replacing it same with x

jabez177 (jabez177):

m is the slope What is the slope in the question?

OpenStudy (tparker22):

The question says -3

pooja195 (pooja195):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @tparker22 Write the point-slope form of the line that passes through (1, -5) and is perpendicular to a line with a slope of 1. \(\color{#0cbb34}{\text{End of Quote}}\) It says "slope of 1" m=slope

OpenStudy (tparker22):

Oops, I'm sorry! I was looking at the wrong question. I clicked on the wrong one, okay I'm back to the one we're all currently on

OpenStudy (tparker22):

So it would really be y - y5 = m(x - x1) ?

OpenStudy (tparker22):

I mean, I can plug in the info, but I can't work it. That's my issue If one of y'all could explain in detail step by step, that'd be perfect and maybe I could take it from there with the rest of my problems

OpenStudy (leenathan):

pooja you stole my qustion! oh well i need to go bye

pooja195 (pooja195):

Basically we have a point slope formula \[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1) In this particular problem you are given that: x=1 y=-5 slope which is m=1 You take this information and you plug it in place of the variables present in the formula

OpenStudy (tparker22):

Okay, I get that so far

pooja195 (pooja195):

Can you tell me, what equation we would have once we plug in our numbers? Keeping in mind that when two minus signs are next to each other it becomes a +

OpenStudy (tparker22):

We would have ----- y - y5 = 1(x-x1) ?

OpenStudy (tparker22):

or would every y be 5 and every x be 1?

pooja195 (pooja195):

You are correct thus with a minor mistake y +5 = 1(x-x1)

pooja195 (pooja195):

Continue with plugging in your x value

OpenStudy (tparker22):

Ok so since we have this - y + 5 = 1(x-x1) - then we would continue with plugging in the x value you said. My current x value is 1, so would be be y + 5 = 1(1-x1)? see, this is where I get confused, I'm sorry :/

pooja195 (pooja195):

You need to plug your x value in the x1 spot

OpenStudy (tparker22):

That would be 1 * 1 then, which is 1?

pooja195 (pooja195):

No x1 is part of the coordinate (1,-5) x1=1

OpenStudy (tparker22):

Ok

pooja195 (pooja195):

plug it in

OpenStudy (tparker22):

Okay, so I plugged these numbers in and did a bit more. Would this be it?

OpenStudy (tparker22):

(y-y1)=m(x-x1) (y--5)=1(x-1) y=1x-1+-5 y=1x-6 1x-y-6=0

pooja195 (pooja195):

(y--5)=1(x-1) Correct but \[\huge~\rm~\bf~ y+5=1(x-1)\] \[\huge~\rm~\bf~ y+5=1x-1\] \[\huge~\rm~\bf~ y=1x-6\] \[Perpendicular~slope=Flipped~sign+Reciprocal~(flipped fraction) \]

OpenStudy (tparker22):

Ok! :)

pooja195 (pooja195):

Whats the opposite of 1?

OpenStudy (tparker22):

-1

OpenStudy (tparker22):

So parallel is positive, perpendicular is opposite?

OpenStudy (tparker22):

-negative I meant

pooja195 (pooja195):

Not really I suggest you use this site to help you differentiate between the 2 https://www.mathsisfun.com/algebra/line-parallel-perpendicular.html

OpenStudy (tparker22):

Or is it the other way around?

OpenStudy (tparker22):

Okay, I'll look into it! Thank you!

pooja195 (pooja195):

Parallel--->slopes stay the same Perpendicular---->Flip the fraction and if there isn't one make one by adding a 1 to the bottom, and flip the sign (positive # goes to negative; negative # goes to positive)

OpenStudy (tparker22):

okay! thank you :)

pooja195 (pooja195):

You're welcome.